M-Structures: Extending a Parallel, Non-strict, Functional Language with State

Computation Structures Group Memo 327 March 18, 1991

Paul S. Barth Rishiyur S. Nikhil Arvind

In Proceedings on Functional Programming and Computer Architecture, Cambridge, MA, August 28-30, 1991.

This report describes research done at the Laboratory for Computer Science of the Massachusetts Institute of Technology. Funding for the Laboratory is provided in part by the Advanced Research Projects Agency of the Department of Defense under the Office of Naval Research contract N00014-89-J-1988.